Grade 11General Physics

Can two forces of magnitude 6N and 10N respectively have a resultant of (a). 12N (b). 18N ?

Anmol chaudhary
9 Years agoGrade 11

1 Answer

Shreya D Nanda
ApprovedApproved Tutor Answer9 Years ago
Magnitude of resultant force is given by,  F_{r} =\sqrt{F_{1}^{2} +F_{2}^{2} +2F_{1}F_{2} \cos \theta }
According to the given question, let us consider
  • 12N as answer 
     \therefore 12 = \sqrt{6^2 + 10^2 + 2*6*10 \cos \theta }      \Rightarrow 12=\sqrt{36+100+120\cos\theta } 
 
       or   12^2=136+120\cos\theta      \Rightarrow \cos\theta = [144-136]/120 = 0.667
       
       \Rightarrow \theta = 86.177^{\circ}
 \therefore The answer to the question is 12N. , i.e,  the resultant of forces 6N and 10N can be 12N when angle between the two forces is   \cos^{-1}0.667   or  86.1774^{\circ}
 
 
    NOTE: The resultant of 6N and 10N can never be 18 N,  as from above method, we get    \cos \theta = 1.556
               which is impossible becoz, -1\leq \cos \theta \leq 1
 
 
Hope you understand....